# Resource Allocation in Distributed Quantum Computing Interconnect Networks

### v1

@@ Abstract
Distributed quantum computing (DQC) has emerged as a promising approach to overcome the scalability limitations of monolithic quantum processors in terms of computing capability. However, realising the full potential of DQC requires effective resource allocation. This involves efficiently distributing quantum circuits across the network by assigning each circuit to an optimal subset of quantum processing units (QPUs), based on factors such as their computational power and connectivity. In heterogeneous DQC networks with arbitrary topologies and non-identical QPUs, resource allocation becomes a complex challenge. This paper addresses the problem of resource allocation in such networks, focusing on computing resource management in a quantum farm setting. We propose a multi-objective optimisation algorithm for optimal QPU allocation that aims to minimise the degradation caused by inter-QPU communication latencies due to qubit decoherence, while maximising the number of concurrently assignable quantum circuits. The algorithm takes into account several key factors, including the network topology, QPU characteristics, and quantum circuit structure, to make efficient allocation decisions. We formulate the optimisation problem as a mixed-integer linear program and solve it using standard optimisation tools. Simulation results demonstrate the effectiveness of the proposed algorithm in minimising communication costs and improving resource utilisation compared to a benchmark greedy allocation approach. To complement our proposed QPU allocation method, we also present a compatible quantum circuit scheduling model. Our work provides valuable insights into resource allocation strategies for DQC systems and contributes to the development of efficient execution management frameworks for quantum computing.
@@ I Introduction
Quantum computing has emerged as a promising solution for tackling intractable problems, due to its capacity to solve them significantly faster than traditional computers. In recent years, there have been notable advancements in quantum hardware and control systems, leading to the development of noisy intermediate-scale quantum processing units (QPUs). However, despite these efforts, current quantum processors still remain limited in their computational power. In this context, distributed quantum computing (DQC) has emerged as a promising approach to address the scalability limitations of monolithic quantum processors in terms of the number of qubits. [1, 2, 3, 4].
@@
One of the key requirements for the distributed execution of quantum algorithms is the ability to perform quantum operations between distant qubits residing on separate QPUs. A notable approach to accomplish such remote quantum operations, termed remote gates, utilises two primitives: cat-entangler and cat-disentangler, as proposed by Yimsiriwattana and Lomonaco [1]. This approach involves the utilisation of an entangled pair, local quantum operations on individual QPUs, and classical communication. The cat-entangler creates an entangled state, called a “cat-like state”, between two or more qubits, while the cat-disentangler performs the reverse operation. Figure 1 illustrates the circuit diagram for implementing a controlled-NOT (CNOT) gate based on this approach between two computing qubits, denoted by $q_{\rm cp1}$ and $q_{\rm cp2}$, belonging to separate QPUs (${\rm QPU}_{1}$ and ${\rm QPU}_{2}$). The initial step in this method involves generating entanglement between two dedicated communication qubits, denoted by $q_{\rm cm1}$ and $q_{\rm cm2}$ in Fig. 1, which facilitate the quantum communication between the QPUs. As depicted in the figure, the X operation on ${\rm QPU}_{2}$ and the Z operation on ${\rm QPU}_{1}$ are conditioned on the measurement outcome transmitted classically from the other QPU. This demonstrates that the overall process of remote gate execution involves both quantum and classical communication between QPUs.
@@
This work focuses on QPU allocation and QCirc scheduling, crucial aspects of execution management in a quantum farm. Quantum computing’s unique characteristics, particularly limited qubit coherence time, necessitate resource allocation algorithms that consider not only resource utilisation but also the impact of communication latencies on computation accuracy. Unlike classical computing, where communication latencies primarily affect task completion time, in quantum computing, they can significantly impact computation accuracy. To mitigate qubit decoherence caused by communication delays, effective QPU allocation algorithms must consider the network topology and characteristics, along with the QPUs’ features and limitations, including their capacity (number of computing qubits) and decoherence properties. Additionally, previous research [7] has shown that quantum circuits exhibit different levels of sensitivity to the errors arising from their distributed execution. Therefore, a QPU allocation method that takes the distinctive features of QCircs into account would enhance the overall performance and promote fairness in terms of computation accuracy. Moreover, as QCirc scheduling and QPU allocation are closely interrelated, an efficient scheduling method tailored to the developed QPU allocation strategy is a prerequisite for optimal performance. These factors serve as the primary motivation for this work.
In the following subsections, related work is first outlined, followed by a detailed description of our contributions.
@@ I-A Related work
In the literature, extensive research has focused on quantum compilation and/or circuit partitioning for individual quantum circuits, while relatively limited research efforts have addressed execution management, including QPU allocation. Given the strong interdependencies among these tasks, we present key related works on these topics in what follows.
@@ I-B Our contributions
Despite the growing body of literature on compilation and circuit partitioning techniques for individual quantum circuits, there is a notable lack of research addressing execution management in DQC interconnect networks, particularly in the context of quantum farms. In a quantum farm, multiple requests for executing quantum algorithms are submitted simultaneously, raising important questions about optimal resource allocation and scheduling. For instance, determining the appropriate number of partitions for each quantum circuit and the specific QPUs that should be allocated to each circuit becomes a critical challenge. This work aims to address these gaps through the following contributions.
@@ II Network Model
We consider a DQC network in which quantum computers, as quantum nodes, are interconnected by quantum and classical links, as illustrated in Fig. 3. We assume that the network may have two types of quantum links: direct and multi-hop. In the case of direct links, two nodes share an entangled pair without the use of any intermediate hop, while in the case of multi-hop links, entanglement swapping is utilised to generate end-to-end entanglement between two nodes not interconnected by a direct quantum link [18]. Regarding the quantum nodes, they may have different capabilities; for instance, some nodes may have entanglement swapping capability, allowing them to function as quantum repeaters (intermediate hops). Furthermore, the QPUs within the nodes may have varying capacities and decoherence properties. These assumptions enable us to consider a highly general quantum network model, where neither the nodes nor the quantum links are required to be identical. Moreover, the logical network topology can be fully-connected or partially-connected. In the following, we provide a more detailed discussion on the quantum and classical links and their associated latencies.
@@ II-B Classical communication latency
As depicted in Fig. 1, inter-QPU classical communication is an essential component of remote gate execution. We denote the classical communication latency between the QPU pair ${\rm{QPU}}{j_{1}}$ and ${\rm{QPU}}{j_{2}}$ as $T^{\rm cl}_{j_{1}j_{2}}$. If direct communication between QPUs can be completed in hardware without transferring to the software domain, lower $T^{\rm cl}_{j_{1}j_{2}}$ values and more deterministic behaviour can be achieved. Programmable hardware such as FPGAs plays a key role in the direct interconnection between nodes. These interconnections can be deterministically established within nanoseconds when FPGAs connect the QPUs via optical links for 10GbE, as reported in [7].
@@ III Scheduling Model
The proposed workflow for the QCircs scheduling is presented in Figure 4. The QCircs queued in the QCircs Queue are assigned a priority using a predefined mechanism that may consider parameters such as dependency to other quantum programs, deadline, and wait time. Although the specifics of the priority assignment are beyond the scope of this paper, the wait time of a QCirc significantly influences its priority, preventing starvation from prolonged queuing. We assume that priority assignment occurs regularly, with the QCircs in the queue sorted accordingly, and that each QCirc has a priority $p$.
@@ IV-A1 Objective 1: minimising the cost imposed by inter-QPU communication
$P_{m}^{(c)}=\prod_{\begin{subarray}{c}j_{1},j_{2},j_{3}\\ j_{2}>j_{1}\end{subarray}}{a_{j_{1}j_{2}j_{3}}^{\quad x_{mj_{3}}n^{(rg)}_{mj_{1}j_{2}}f_{mj_{1}}f_{mj_{2}}}}$ (9)
@@
We define the cost imposed by inter-QPU communication, corresponding to the matrix $X$, as follows:
$cost_{m}(X)\triangleq P_{m}^{(e)}=1-P_{m}^{(c)}$ (10)
The above cost function represents how latencies from quantum and classical links used in remote gate executions could degrade the accuracy of a quantum computing task. While this error probability metric does not directly quantify post-measurement errors in a quantum circuit’s output, it establishes an upper bound on these errors when focusing exclusively on latency-induced noise, without considering other noise sources. In the remainder of this subsection, we will derive a simplified expression for this cost function.
In the first step, the above cost function can be simplified by utilising the function $h(y)=|\log_{2}{(1-y)}|$, which exhibits strict monotonicity and injectiveness within the interval $[0,1)$, as follows:
$C_{m}^{\text{s1}}(X)=h(cost_{m}(X))=\quad\quad\quad\quad\quad$
@@ V Evaluation and results
This section evaluates the proposed QPU allocation algorithm through simulations. We consider a DQC network with 20 QPUs. To investigate the algorithm’s performance under varying network topologies, we examine two primary topology types: a grid topology with QPUs arranged in a 5x4 grid, and random topologies based on Erdős-Rényi graph model [24]. The latter are generated using NetworkX [25], where each edge is generated with a probability $p_{\rm ed}$. We consider four cases for $p_{\rm ed}$: 0.1, 0.2, 0.5, and 1. Note that $p_{\rm ed}=1$ corresponds to a fully-connected topology. To ensure the random topologies are connected graphs, the graph generation process is repeated until a connected one is achieved. These diverse topologies allow us to assess the algorithm’s performance across a range of network structures, from regular grids to varying degrees of random connectivity.
In our simulation, we assume the quantum network supports entanglement swapping, enabling the generation of end-to-end entanglement between non-adjacent nodes. For the latency parameters, $T^{\rm eg}_{j_{1}j_{2}}$ and $T^{\rm cl}_{j_{1}j_{2}}$, we assume a simple model where the route between any two nodes ${\rm{QPU}}_{j_{1}}$ and ${\rm{QPU}}_{j_{2}}$ is the shortest path, and the values of $T^{\rm eg}_{j_{1}j_{2}}$ and $T^{\rm cl}_{j_{1}j_{2}}$ are given by
$T^{\rm cl}_{j_{1}j_{2}}=T^{\rm cl}_{{\rm el}}(n_{\rm s}+1),$ (18)
To evaluate the effectiveness of our proposed algorithm, we adopt a greedy allocation approach as the benchmark, rather than a random allocation baseline. The benchmark algorithm prioritises QPU capacity utilisation, making allocation decisions solely based on QPU capacity and QCirc width. It iteratively matches each QCirc to an appropriate QPU over a series of rounds, aiming to minimise the gap between the QCirc’s required qubits and the available capacity of the allocated QPU in each round. The benchmark method, described in Algorithm 1, can be adapted to both network scenarios (a) and (b) by setting the parameter $R$ to a large number or one, respectively.
We evaluate four main metrics, averaged over 100 iterations:
a) The ratio of quantum circuits successfully assigned to the total $M$ QCircs in the batch, denoted by $r_{\rm asg}$.
b) The inter-QPU communication cost $C_{\rm com}$ (equation (3)).
c) The total number of sub-circuits resulting from partitioning, denoted by $N_{\rm part}$.
d) The total number of remote gates resulting from circuit partitioning, denoted by $N_{\rm rgate}$.
To account for the varying number of successfully assigned circuits, the average values of $C_{\rm com}$, $N_{\rm part}$, and $N_{\rm rgate}$ are normalised by average $r_{\rm asg}$.
Our simulation results demonstrate that for the Unlimited R scenario, all QCircs within a batch are successfully assigned, resulting in an assignment ratio ($r_{\rm asg}$) of 1 in all examined cases, encompassing both the proposed and benchmark methods. This outcome aligns with our expectations, given that the average qubit requirement of all QCircs matches the average total number of available qubits, and the condition $R=\infty$ imposes no restrictions on QCirc allocation. However, this behavior changes when $R$ is set to 1 in Limited R scenario, as detailed in Table II. In all instances, $r_{\rm asg}$ is less than one, although the proposed method consistently outperforms the benchmark in terms of assignment ratio.
TABLE II: Assignment ratio, $r_{\rm asg}$, for the Limited R scenario.
$M=14$ 0.917 0.917 0.917 0.917 0.917 0.806
$M=10$ 0.975 0.96 0.987 0.977 0.974 0.912
$M=8$ 0.9838 0.995 0.9863 0.995 0.995 0.896
Figure 5 presents the normalised inter-QPU communication cost, $C_{\rm com}$, for all three QCirc configurations examined in our simulations. The results clearly demonstrate that the proposed method substantially reduces the inter-QPU communication cost, $C_{\rm com}$, compared to the benchmark approach. Moreover, by comparing results across different edge probabilities in random topologies, it becomes evident that this improvement is more pronounced in less connected network structures. The benchmark method exhibits a strong sensitivity to network topology, with $C_{\rm com}$ consistently increasing as edge probability, $p_{\rm ed}$, decreases. In contrast, the proposed method is significantly less affected. Although a slight decrease in $C_{\rm com}$ is observed in certain cases, particularly for $M=8$, this reduction is considerably less pronounced than that of the benchmark method. This is primarily because the benchmark method does not consider the network topology and characteristics in its allocation decisions.
Figures 6(a) and (b) depict the normalised $N_{\rm rgate}$ and $N_{\rm part}$, respectively, for all scenarios considered in our simulations. These figures demonstrate that the proposed method effectively reduces the number of circuit partitions and remote gate operations, compared to the benchmark method. These improvements can be attributed to several key factors. Firstly, the proposed method accounts for network topology, link latencies, QPU capacities, and QCirc characteristics, while the benchmark method solely considers QPU capacities and circuit widths. Another critical advantage is the proposed method’s holistic approach, where allocation decisions are made by considering the entire batch of QCircs, in contrast to the benchmark method’s one-by-one assignment approach. The latter can lead to inefficient resource allocation and excessive partitioning, especially for QCircs assigned later in the queue, when a significant portion of the computing resources has already been allocated. This can potentially result in suboptimal partitioning and an increased number of remote gates, particularly for QCircs with high connectivity (corresponding to higher $\nu_{m}$). The proposed method’s batch-wise allocation strategy mitigates these issues, leading to more efficient resource utilisation, reduced partitioning, and fewer remote gate operations.
To examine more closely the impact of inter-QPU communication latencies, we assess the error probability, $P^{(e)}$ (defined in Equation (10)), for each QCirc type (GHZ, VQE, DJ, and QFT). In our simulations, this metric is averaged across all instances of a given circuit type over 100 simulation runs. Our primary focus is on random topologies with varying edge probabilities, $p_{\rm ed}$, under the Limited R scenario. Figure 7 illustrates the average $P^{(e)}$ as a function of $p_{\rm ed}$ for different circuit configurations and types. Notably, GHZ circuits exhibit the lowest latency-induced errors, whereas QFT circuits generally have the highest, across most scenarios. This correlation is attributed to the connectivity patterns of these circuits. As shown in Fig. 8 for QCircs of width $w=4$, GHZ has the sparsest connectivity, whereas QFT is fully connected.
Algorithm 1 Benchmark QPU allocation algorithm
@@
Figure 5: Comparison of proposed and benchmark QPU allocation methods based on the normalised $C_{\rm com}$ across various network scenarios and topologies. Three main cases are considered for the value of $M$ and the width of each QCirc: 1) $M=14$, $w_{m}\in\{15,\ldots,25\}$, 2) $M=10$, $w_{m}\in\{23,\ldots,33\}$, and 3) $M=8$, $w_{m}\in\{30,\ldots,40\}$.
Figure 6: Performance comparison of the proposed and benchmark QPU allocation methods across various network scenarios and topologies. Three main cases are considered for the value of $M$ and the width of each QCirc: 1) $M=14$, $w_{m}\in\{15,\ldots,25\}$, 2) $M=10$, $w_{m}\in\{23,\ldots,33\}$, and 3) $M=8$, $w_{m}\in\{30,\ldots,40\}$. (a) Normalised $N_{\rm rgate}$. (b) Normalised $N_{\rm part}$.
Figure 7: Error probability as a function of edge probability for random graph topologies and different quantum circuit types, assuming single circuit partition per QPU ($R=1$). Three main cases are considered for the value of $M$ and the width of each QCirc: 1) $M=14$, $w_{m}\in\{15,\ldots,25\}$, 2) $M=10$, $w_{m}\in\{23,\ldots,33\}$, and 3) $M=8$, $w_{m}\in\{30,\ldots,40\}$.
@@
Figure 9: Comparison of proposed and benchmark QPU allocation methods, considering the specific case where the number of partitions is limited to one for the QFT circuit. A scenario with Limited R, $M=14$, and Grid topology is assumed. (a) Normalised $C_{\rm com}$ (b) Normalised $N_{\rm part}$ (b) Normalised $N_{\rm rgate}$.
@@
Lastly, we evaluate the performance of the proposed method when the number of partitions for highly-connected QCircs is restricted. To this end, we consider the specific case where the number of partitions is limited to one ($K^{\rm max}=1$) for the QFT circuit, which is the QCirc with the largest connectivity parameter, $\nu$, among the four quantum circuits used in our simulations. The remaining three QCircs have no restrictions on their partition count. We focus on the scenario of Limited R, $M=14$, and Grid topology. For a fair comparison, the benchmark method is modified to accommodate this limitation. Specifically, QFT circuits are only assigned to QPUs with capacities exceeding their circuit widths. This modified version is referred to as “M-benchmark”.

### v2

# Resource Management and Circuit Scheduling for Distributed Quantum Computing Interconnect Networks

@@ Abstract
Distributed quantum computing (DQC) has emerged as a promising approach to overcome the scalability limitations of monolithic quantum processors in terms of computing capability. However, realising the full potential of DQC requires effective resource allocation. This involves efficiently distributing quantum circuits across the network by assigning each circuit to an optimal subset of quantum processing units (QPUs), based on factors such as their computational power and connectivity. In heterogeneous DQC networks with arbitrary topologies and non-identical QPUs, resource allocation becomes a complex challenge. This paper addresses the problem of resource allocation in such networks, focusing on computing resource management in a quantum farm setting. We propose a multi-objective optimisation algorithm for optimal QPU allocation that aims to minimise the degradation caused by inter-QPU communication latencies due to qubit decoherence, while maximising the number of concurrently assignable quantum circuits. The algorithm takes into account several key factors, including the network topology, QPU characteristics, and quantum circuit structure, to make efficient allocation decisions. We employ mixed integer linear programming to solve this optimisation problem. Simulation results demonstrate the effectiveness of the proposed algorithm in minimising communication costs and improving resource utilisation compared to a benchmark greedy allocation approach. Notably, assuming a single circuit partition per QPU, the success rate of quantum circuit assignments improves by 5.25%-13.75%. To complement our proposed QPU allocation method, we also present a compatible quantum circuit scheduling model. Our work provides valuable insights into resource allocation strategies for DQC systems and contributes to the development of efficient execution management frameworks for quantum computing.
@@ I Introduction
Quantum computing has gained attention as a solution for tackling intractable problems, due to its capacity to solve them significantly faster than traditional computers. In recent years, there have been notable advancements in quantum hardware and control systems, leading to the development of noisy intermediate-scale quantum processing units (QPUs). However, despite these efforts, current quantum processors still remain limited in their computational power. In this context, distributed quantum computing (DQC) can offer a substantial benefit in scaling up quantum computing beyond the constraints of monolithic systems [1, 2, 3, 4].
@@
One of the key requirements for the distributed execution of quantum algorithms is the ability to perform quantum operations between distant qubits residing on separate QPUs. A notable approach to accomplish such remote quantum operations, termed remote gates, utilises two primitives: cat-entangler and cat-disentangler, as proposed by Yimsiriwattana and Lomonaco [1]. This method involves the utilisation of an entangled pair, local quantum operations on individual QPUs, and classical communication. Consequently, the overall process of remote gate execution incorporates both quantum and classical communication between QPUs.
@@
This work focuses on QPU allocation and QCirc scheduling, crucial aspects of execution management in a quantum farm. In the following subsections, related work is first outlined, followed by a detailed description of our contributions.
@@ I-A Related work
In the literature, extensive research has focused on quantum compilation and circuit partitioning, with fewer efforts on execution management. Given the strong interdependencies among these tasks, we present key related works on these topics in what follows.
@@ I-B Our contributions
Despite the growing body of literature on compilation and circuit partitioning for individual quantum circuits, there is a notable lack of research addressing execution management in DQC networks, particularly in the context of quantum farms. In a quantum farm, multiple quantum algorithm requests are submitted simultaneously, raising key challenges in resource allocation and scheduling, such as determining the number of partitions and the optimal QPUs for each quantum circuit.
Quantum computing’s unique constraints, like limited qubit coherence time, demand resource management strategies that account for both resource utilisation and communication latencies. Unlike classical computing, where latency mainly affects timing, it can degrade computation accuracy due to qubit decoherence in quantum computing. Furthermore, previous research [17] has shown that quantum circuits exhibit varying levels of sensitivity to the errors arising from their distributed execution. Therefore, a QPU allocation method that accounts for network topology, QPU characteristics, and distinctive features of QCircs can enhance the overall performance and promote fairness in computation accuracy. In the following, we briefly outline our contributions to addressing these challenges.
@@ II Network Model
We consider a DQC network in which quantum computers, as quantum nodes, are interconnected by quantum and classical links, as illustrated in Fig. 2. We assume that the network may have two types of quantum links: direct and multi-hop. In the case of direct links, two nodes share an entangled pair without the use of any intermediate hop, while in the case of multi-hop links, entanglement swapping is utilised to generate end-to-end entanglement between two nodes not interconnected by a direct quantum link [18]. Regarding the quantum nodes, they may have different capabilities; for instance, some nodes may have entanglement swapping capability, allowing them to function as quantum repeaters (intermediate hops). Furthermore, the QPUs within the nodes may have varying capacities and decoherence properties. These assumptions enable us to consider a highly general quantum network model, where neither the nodes nor the quantum links are required to be identical. Moreover, the logical network topology can be fully-connected or partially-connected.
We assume that the DQC network utilises the method proposed in [1] to perform a remote gate. Figure 3 illustrates the circuit diagram for implementing a controlled-NOT (CNOT) gate based on this approach between two computing qubits, denoted by $q_{\rm cp1}$ and $q_{\rm cp2}$, belonging to separate QPUs (${\rm QPU}_{1}$ and ${\rm QPU}_{2}$). The initial step in this method involves generating entanglement between two dedicated communication qubits, denoted by $q_{\rm cm1}$ and $q_{\rm cm2}$ in Fig. 3, which facilitate the quantum communication between the QPUs. As depicted in the figure, the X operation on ${\rm QPU}_{2}$ and the Z operation on ${\rm QPU}_{1}$ are conditioned on the measurement outcome transmitted classically from the other QPU. In the following, we provide a more detailed discussion on the quantum and classical links within the DQC network, which are essential for remote gate execution, along with their associated latencies.
@@ II-B Classical communication latency
As depicted in Fig. 3, inter-QPU classical communication is an essential component of remote gate execution. We denote the classical communication latency between the QPU pair ${\rm{QPU}}{j_{1}}$ and ${\rm{QPU}}{j_{2}}$ as $T^{\rm cl}_{j_{1}j_{2}}$. If direct communication between QPUs can be completed in hardware without transferring to the software domain (such as IP), lower $T^{\rm cl}_{j_{1}j_{2}}$ values and more deterministic behaviour can be achieved. Programmable hardware such as FPGAs plays a key role in the direct interconnection between nodes. These interconnections can be deterministically established within nanoseconds when FPGAs connect the QPUs via optical links for 10GbE, as reported in [17].
@@ III Scheduling Model
The proposed workflow for the QCircs scheduling is presented in Figure 4. The QCircs queued in the QCircs Queue are assigned a priority using a predefined mechanism that may consider parameters such as dependency to other quantum programs and wait time. Although the specifics of the priority assignment are beyond the scope of this paper, the wait time of a QCirc significantly influences its priority, preventing starvation from prolonged queuing. We assume that priority assignment occurs regularly, with the QCircs in the queue sorted accordingly, and that each QCirc has a priority $p$.
@@ IV-A1 Objective 1: minimising the cost imposed by inter-QPU communication
$P_{m}^{(c)}(X)=\prod_{\begin{subarray}{c}j_{1},j_{2},j_{3}\\ j_{2}>j_{1}\end{subarray}}{a_{j_{1}j_{2}j_{3}}^{\quad x_{mj_{3}}n^{(rg)}_{mj_{1}j_{2}}f_{mj_{1}}f_{mj_{2}}}}$ (9)
@@
We define the error probability imposed by inter-QPU communication, corresponding to the matrix $X$, as follows:
$P_{m}^{(e)}(X)=1-P_{m}^{(c)}(X)$ (10)
The above error probability function represents how latencies from quantum and classical links used in remote gate executions could degrade the accuracy of a quantum computing task. While this error probability metric does not directly quantify post-measurement errors in a quantum circuit’s output, it establishes an upper bound on these errors when focusing exclusively on latency-induced noise, without considering other noise sources. Since the performance of the QPU allocation algorithm is highly associated to the error probability, we define the cost function taking the error probability into account utilising the function $h(y)=|\log_{2}{(1-y)}|$ as follows:
$C_{m}^{\text{s1}}(X)=h(P_{m}^{(e)}(X))=\quad\quad\quad\quad\quad$
@@ V Evaluation and results
This section evaluates the proposed QPU allocation algorithm through simulations. We consider a DQC network with 20 QPUs. To investigate the algorithm’s performance under varying network topologies, we examine two primary topology types: a grid topology with QPUs arranged in a 5x4 grid, and random topologies based on Erdős-Rényi graph model [24]. The latter are generated using NetworkX [25], where each edge is generated with a probability $p_{\rm ed}$. We consider four cases for the edge probability, $p_{\rm ed}$: 0.1, 0.2, 0.5, and 1. Note that $p_{\rm ed}=1$ corresponds to a fully-connected topology. To ensure the random topologies are connected graphs, the graph generation process is repeated until a connected one is achieved. These diverse topologies allow us to assess the algorithm’s performance across a range of network structures, from regular grids to varying degrees of random connectivity.
Benchmark Algorithm Greedy QPU allocation
@@
TABLE II: Assignment ratio, $r_{\rm asg}$, for the Limited R scenario.
$M=14$ 0.917 0.917 0.917 0.917 0.917 0.806
$M=10$ 0.975 0.96 0.987 0.977 0.974 0.912
$M=8$ 0.9838 0.995 0.9863 0.995 0.995 0.896
In our simulation, we assume all DQC network nodes can act as quantum repeaters, allowing end-to-end entanglement between any pair. For the latency parameters, $T^{\rm eg}_{j_{1}j_{2}}$ and $T^{\rm cl}_{j_{1}j_{2}}$, we assume a simple model where the route between any two nodes ${\rm{QPU}}_{j_{1}}$ and ${\rm{QPU}}_{j_{2}}$ is the shortest path, and the values of $T^{\rm eg}_{j_{1}j_{2}}$ and $T^{\rm cl}_{j_{1}j_{2}}$ are given by
$T^{\rm cl}_{j_{1}j_{2}}=T^{\rm cl}_{{\rm el}}(n_{\rm s}+1),$ (18)
Figure 5: Performance comparison of the proposed and benchmark QPU allocation methods based on the scaled communication cost ($C_{\rm com}^{(s)}$) across various network scenarios and topologies. Three main cases are considered based on the number of QCircs ($M$) and the width of each QCirc ($w_{m}$): 1) $M=14$, $w_{m}\in\{15,\ldots,25\}$, 2) $M=10$, $w_{m}\in\{23,\ldots,33\}$, and 3) $M=8$, $w_{m}\in\{30,\ldots,40\}$.
Figure 6: Performance comparison of the proposed and benchmark QPU allocation methods across various network scenarios and topologies. Three main cases are considered based on the number of QCircs ($M$) and the width of each QCirc ($w_{m}$): 1) $M=14$, $w_{m}\in\{15,\ldots,25\}$, 2) $M=10$, $w_{m}\in\{23,\ldots,33\}$, and 3) $M=8$, $w_{m}\in\{30,\ldots,40\}$. The metrics evaluated are: (a) Scaled remote gate count ($N_{\rm rgate}^{(s)}$), and (b) Scaled partition count ($N_{\rm part}^{(s)}$).
Figure 7: Error probability ($P^{(e)}$) for different QCirc types as a function of the edge probability parameter ($p_{\rm ed}$), which indicates network connectivity for random topologies. A single circuit partition per QPU ($R=1$) is assumed. Three main cases are considered based on the number of QCircs ($M$) and the width of each QCirc ($w_{m}$): 1) $M=14$, $w_{m}\in\{15,\ldots,25\}$, 2) $M=10$, $w_{m}\in\{23,\ldots,33\}$, and 3) $M=8$, $w_{m}\in\{30,\ldots,40\}$.
@@
To evaluate the effectiveness of our proposed algorithm, we adopt a greedy allocation approach as the benchmark, rather than a random allocation baseline. The benchmark algorithm prioritises QPU capacity utilisation, making allocation decisions solely based on QPU capacity and QCirc width. It iteratively matches each QCirc to an appropriate QPU over a series of rounds, aiming to minimise the gap between the QCirc’s required qubits and the available capacity of the allocated QPU in each round. The benchmark method, described in Benchmark Algorithm, can be adapted to both network scenarios (a) and (b) by setting the parameter $R$ to a large number or one, respectively.
We define and evaluate four main metrics:
a) Assignment ratio ($r_{\rm asg}$): The ratio of successfully assigned QCircs to the total $M$ QCircs in the batch, averaged over 100 iterations.
b) Scaled communication cost ($C_{\rm com}^{(s)}$): The inter-QPU communication cost $C_{\rm com}$ (equation (3)), averaged over 100 iterations and scaled by the assignment ratio to provide a fair comparison.
c) Scaled partition count ($N_{\rm part}^{(s)}$): The total number of sub-circuits resulting from partitioning of successfully assigned QCircs, averaged over 100 iterations and scaled by the assignment ratio.
d) Scaled remote gate count ($N_{\rm rgate}^{(s)}$): The total number of remote gates resulting from partitioning of successfully assigned QCircs, averaged over 100 iterations and scaled by the assignment ratio.
Our simulation results demonstrate that for the Unlimited R (single circuit partition per QPU) scenario, all QCircs within a batch are successfully assigned, resulting in an assignment ratio ($r_{\rm asg}$) of 1 in all examined cases, encompassing both the proposed and benchmark methods. This outcome aligns with our expectations, given that the average qubit requirement of all QCircs matches the average total number of available qubits, and the condition $R=\infty$ imposes no restrictions on QCirc allocation. However, this behavior changes when $R$ is set to 1 in Limited R scenario, as detailed in Table II. In all instances, $r_{\rm asg}$ is less than one, with the proposed method consistently outperforming the benchmark. These results demonstrate better resource utilisation for the proposed method.
Figure 5 presents the scaled communication cost, $C_{\rm com}^{(s)}$, for all three QCirc configurations examined in our simulations. The results clearly demonstrate that the proposed method substantially reduces the inter-QPU communication error cost compared to the benchmark approach. Moreover, by comparing results across different edge probabilities in random topologies, it becomes evident that this improvement is more pronounced in less connected network structures. The benchmark method exhibits a strong sensitivity to network topology. As edge probability ($p_{\rm ed}$) decreases, which corresponds to a reduced degree of connectivity in the network topology, $C_{\rm com}^{(s)}$ consistently increases. In contrast, the proposed method is significantly less affected. Although a slight decrease in $C_{\rm com}$ is observed in certain cases, particularly for $M=8$, this reduction is considerably less pronounced than that of the benchmark method. This is primarily because the benchmark method does not consider the network topology and characteristics in its allocation decisions.
Figures 6 (a) and (b) depict the scaled remote gate count ($N_{\rm rgate}^{(s)}$) and scaled partition count ($N_{\rm part}^{(s)}$) for all scenarios considered in our simulations. These figures demonstrate that the proposed method effectively reduces the number of circuit partitions and remote gate operations, compared to the benchmark method. These improvements can be attributed to several key factors. Firstly, the proposed method accounts for network topology, link latencies, QPU capacities, and QCirc characteristics, while the benchmark method solely considers QPU capacities and circuit widths. Another critical advantage is the proposed method’s holistic approach, where allocation decisions are made by considering the entire batch of QCircs, in contrast to the benchmark method’s one-by-one assignment approach. The latter can lead to inefficient resource allocation and excessive partitioning, especially for QCircs assigned later in the queue, when a significant portion of the computing resources has already been allocated. This can potentially result in suboptimal partitioning and an increased number of remote gates, particularly for QCircs with high connectivity. The proposed method’s batch-wise allocation strategy mitigates these issues, leading to more efficient resource utilisation, reduced partitioning, and fewer remote gate operations.
To examine more closely the impact of inter-QPU communication latencies, we assess the error probability, $P^{(e)}$ (defined in Equation (10)), for each QCirc type (GHZ, VQE, DJ, and QFT). In our simulations, this metric is averaged across all instances of a given circuit type, aggregated from all 100 simulation runs. Our primary focus is on random topologies with varying edge probabilities, $p_{\rm ed}$, under the Limited R (single circuit partition per QPU) scenario. Figure 7 illustrates the average error probability ($P^{(e)}$) as a function of edge probability ($p_{\rm ed}$) for different circuit configurations and types. Notably, GHZ circuits exhibit the lowest latency-induced errors, whereas QFT circuits generally have the highest, across most scenarios. This correlation is attributed to the connectivity patterns of these circuits. As shown in Fig. 8 for QCircs of width $w=4$, GHZ has the sparsest connectivity, whereas QFT is fully connected.
@@
Figure 9: Performance comparison of the proposed and benchmark QPU allocation methods, considering two cases: one where the partition count for the QFT circuit is limited to one ($K^{\rm max}=1$), and another where there is no limitation on $K^{\rm max}$. A scenario with Limited R, $M=14$, and Grid topology is assumed. “M-benchmark” refers to a modified version of the benchmark algorithm that accommodates the limitation on $K^{\rm max}$. The metrics evaluated are: (a) Scaled communication cost ($C_{\rm com}^{(s)}$), (b) Scaled partition count ($N_{\rm part}^{(s)}$), and (c) Scaled remote gate count ($N_{\rm rgate}^{(s)}$).
@@
Lastly, we evaluate the performance of the proposed method when the number of partitions for highly-connected QCircs is restricted. To this end, we consider the specific case where the number of partitions is limited to one ($K^{\rm max}=1$) for the QFT circuit, which is the QCirc with the largest connectivity parameter, $\nu$, among the four quantum circuits used in our simulations. The remaining three QCircs have no restrictions on their partition count. We focus on the scenario of Grid topology, Limited R (single circuit partition per QPU), and number of QCircs equal to 14 ($M=14$) with QCirc width ($w_{m}$) in the range $\{15,\ldots,25\}$. For a fair comparison, the benchmark method is modified to accommodate this limitation. Specifically, QFT circuits are only assigned to QPUs with capacities exceeding their circuit widths. This modified version is referred to as “M-benchmark”.